OpenStudy (anonymous):

solve for g 5g-2h=3g+1 answer in this format g=(__)/__

OpenStudy (paki):

combine the like terms... and leave the "g" term alone...

OpenStudy (anonymous):

g=h+1/2?

OpenStudy (paki):

2h+1/2=g... you have calculated "h" wrong....

OpenStudy (anonymous):

5g-2h=3g+1 add 2h 5g = 3g + 1 + 2h subtract 3g 2g = 1+2h divide by 2 g = (1+2h)/2

OpenStudy (anonymous):

my mistake lol but i got it